# Thermodynamic & pattern-matching efficiency in heterogeneous networks

### v2

@@ 4.2 Performance at criticality
As previously discussed, the $\tau$-scale where the bifurcation in panels (e)-(i) of Figure 4 and in panels (a)-(c) of Figure 5 shows up indicates the scale involved in the structural transition and it is the resolution scale where the thermodynamic efficiency overlaps with the pattern-matching efficiency as a function of $r$. Nevertheless, the critical value of $r$ where such structural transition appears is not necessarily the value where the efficiencies have the global maximum.
@@
To discuss this empirical result we turn our attention to the panels (a)-(c) in Figure 4 and panels (a)-(c) in Figure 5. While for the FashionMNIST dataset there is a rough correspondence between the $r\sim 0.30$ where the structural transition emerges and the $r$-range $r\in(0.30,0.50)$ where the global maximum of the efficiencies is observed, for the MNIST and NIST datasets the thermodynamic and algorithmic efficiencies show a maximum in ranges of $r$ which are farther away from the respective bifurcation points $r\sim 0.35$ and $r\sim 0.45$. The term ”far away” is quantified by the purple-shaded areas Figure 4. We rather observe in panels (b), (c), (e) and (f) of Figure 4 that the maximum efficiencies are posited where the $\tau^{(2)}_{peak}$ have a discontinuous jump in panels (b), (c) of Figure 5 at $r\sim 0.60$ for the MNIST and at $r\sim 0.70$ for the NIST. The argument holds also for the MNIST if we include $\theta$ points in the one standard deviation illustrated by the shaded-purple area. The jump for the MNIST curve in Figure 5 is such that the two $\tau^{(1,2)}_{peak}$ curves overlap. We mention that this effect is mitigated for larger datasets, i.e. larger networks, as we show in Section S2 of the Supplementary. The jump is caused by the presence of a third large peak in between $\tau^{(1)}$ and $\tau^{(2)}$ (see Figure 4 and the entropy curves depicted in Figure S6). Moreover, for the MNIST case, the jump is sufficiently strong to be reflected on a local maximum in the efficiency, $\eta$. Even though this local maximum in $\eta$ is not registered by the averaged pattern-matching efficiency $\theta$ in Figure 4(b), it is reflected in its standard deviation which allows us to include the $r$-range indicated by the shaded purple area, which corresponds to the $\eta$-peak-to-peak $r$-range in Figure 4(e).
Nevertheless, we stress once again that we did not compute the thermodynamic efficiency at the corresponding $\tau$ scale where discontinuity of the $\tau^{(2)}_{peak}$ curve appears, but we rather measured $\eta$ at the $\tau$ scale indicated by the bifurcation.
@@
To explain the diversity in the positioning of the maximum of the efficiencies, either in the vicinity of the transition point or father away, we turn to the argument, already present in the literature [28, 10, 29], that indicates the necessity of tuning a system, or of a system to self-tune, at the border of a phase transition only weather the task requires a high level of complexity. While for the FashionMNIST, a notoriously more complex dataset when compared to both the MNIST and NIST, the optimal value for the efficiency is posited in the proximity of the critical value of $r\sim 0.30$, the optimal values of $r$ for the NIST and MNIST datasets are evidently far away from their respective transition values. The MNIST, which is the intermediate dataset in terms of complexity, is an interesting case as it shows two maxima for the thermodynamic efficiency $\eta$. While the global maximum in Figure 4(e) is far away from the critical value of $r$ illustrated by the bifurcation in panel (b) of Figure 5, the local maxima is at $r\sim 0.35-0.40$ which corresponds to the critical value of $r$ indicated by the bifurcation. This observation is in line with the hypothesis that the system takes advantage from the critical behaviour only when a fairly complex task is presented to it and in line with other observations discussed in the literature [28, 10]. Thus, it stresses the importance of dataset under analyses.
@@ 5 Discussion
When confronted to real world environment, in order to survive, living systems not only need the capability of interpreting the environment, but they also need the capability to quickly react with a coherent response to stimuli possibly drawn from the same distribution they adapted to [33]. Moreover the same system needs to cope with cost of sending information across its structure which can cause an enhanced exploitation of resources. In the context of machine-learning, cost can be quantified by the number of computations to be done before the satisfying answer to the task can be reported, e.g. class-matching. By using the recently developed theoretical tools of statistical physics of fields for networks [5, 4, 34, 21, 3, 2], we have quantified the diversity in the traits of the network with the von Neumann entropy and the transmission flow of information with the Helmholtz free energy. As fastness of information flow and sufficiently rich representation of the outer world are two competing capabilities, the formation process of a complex networked structure is interpreted as the macroscopic thermodynamic outcome of the delicate balance between the two, and a variational principle in terms of thermodynamic efficiency can be devised to explain why a certain structure emerge in place of an other [2]. We have shown that such a thermodynamic efficiency strikingly overlaps with the pattern-matching efficiency defined in terms of accuracy and computational costs, suggesting that the two can be regarded as the same thing but viewed from different angles.
@@
Notably, we have shown that the thermodynamic efficiency needs to be computed at a specific resolution scale $\tau$. We have used the specific heat $C=-dS/d\log\tau$ to identify the $\tau$-resolution which resolves the transition from the trivial-structure, that is a single specific heat peak, to emergence of more complex structures, where the specific heat shows some plateau reflecting the growth of a scale-invariant region where information is integrated at a constant rate. Such $\tau$-resolution scale is akin to an ”ultra-violet” cut-off delimiting the frequency of the network structure across the heterogeneity-transition from star graph to 1d-lattice trough amorphous. While the transition point reflects the network resolution scale where the thermodynamic efficiency needs to be measured, the corresponding transition value of the control parameter $r$ does not coincide with the global maximum in efficiency. We explain this with the statement that a system benefits from the critical point only when it confronts with environments of higher complexity, while it does not benefit from it when the environmental complexity is lower. This is elucidated by the use of three different datasets of increasing complexity and supported by other works in the literature [28, 10, 29].

### v3

@@ 5.1 Performance at the emergence of scale-invariance
As previously discussed in Section 4, in panels (i)-(l) of Figure 6 and in panels (a)-(d) of Figure 7, the $\tau$-scale where the bifurcation shows up indicates the scale involved in the structural transition, and it is the resolution scale where the thermodynamic efficiency for some of the datasets displays consistency with the pattern-matching efficiency as a function of $r$. Nevertheless, the critical value of $r$ where such structural transition appears is not necessarily the value where the efficiencies have the global maximum.
@@
To discuss this empirical result we turn our attention to the panels (a)-(d) and (i)-(l) in Figure 6 and panels (a)-(d) in Figure 7. For the CIFAR10 and FashionMNIST datasets there is a rough correspondence between the $r$-value where the structural transition emerges, $r\sim 0.5$ for CIFAR10 and $r\sim 0.35$ for FashionMNIST, and the $r$-range where the global maximum of the efficiencies is observed, $r\in(0.30,0.50)$ and $r\in(0.35,0.55)$ respectively for the two datasets. On the other hand, for the MNIST and NIST datasets the thermodynamic and pattern-matching efficiencies show a maximum in ranges of $r$ which are farther away from the respective bifurcation points $r\sim 0.4$ and $r\sim 0.5$. The term "far away" is quantified by the purple-shaded areas in panels (a)-(d) of Figure 6. The argument also holds for the MNIST if we include $\theta$ points in the one standard deviation illustrated by the shaded-purple area. For the MNIST case, we observe that a third large peak emerges in between $\tau^{(1)}$ and $\tau^{(2)}$ (see panel (k) in Figure 6 and the entropy curves depicted in Figure S6). This effect is only mitigated for larger datasets, i.e. larger networks, as we show in Section S2 of the Supplementary. This intermediate peak is sufficiently strong to be reflected on a local maximum in the efficiency, $\eta$. Even though this local maximum in $\eta$ is not registered by the averaged pattern-matching efficiency $\theta$ in Figure 6(c), it is reflected in its standard deviation which allows us to include the $r$-range indicated by the shaded purple area, which corresponds to the $\eta$-peak-to-peak $r$-range in Figure 6(g). We stress that also in this case, we measured $\eta$ at the $\tau$ scale indicated by the bifurcation $\tau\sim 10$.
@@
To explain the diversity in the positioning of the maximum of the efficiencies, either in the vicinity of the transition point or farther away, we turn to the argument, already present in the literature [37, 10, 38], that indicates the necessity of tuning a system, or of a system to self-tune, at the border of a phase transition only if a high level of complexity is required by the task, see Figure 2. While for the CIFAR10 and FashionMNIST, more complex datasets when compared to both the MNIST and NIST, their respective optimal value for the efficiency is posited in the proximity of the phase transition, the optimal values of $r$ for the NIST and MNIST datasets are far away from their respective transition values. The MNIST, that is the intermediate dataset in terms of complexity, is an interesting case as it shows two maxima for the thermodynamic efficiency $\eta$. While the global maximum in Figure 6(g) is far away from the critical value of $r$ illustrated by the bifurcation in panel (b) of Figure 7, the local maximum is at $r\sim 0.35-0.40$ which corresponds to the critical value of $r$ indicated by the bifurcation. This observation is in line with the hypothesis that the system takes advantage of the critical behavior of a certain resolution region only when a fairly complex task is presented to it and in line with other observations discussed in the literature [37, 10]. Thus, it stresses the importance of the dataset under analysis and its disordered distribution of points in a high-dimensional feature space.
@@ 6 Discussion
When confronted to real world environment, in order to survive, living systems not only need the capability of interpreting the environment, but they also need the capability to quickly react with a coherent response to stimuli possibly drawn from the same distribution they adapted to [42]. Moreover the same system needs to cope with cost of sending information across its structure which can cause an enhanced exploitation of resources. In the context of machine-learning, cost can be quantified by the number of computations to be done before the satisfying answer to the task can be reported, e.g. class-matching.
The recently developed theoretical tools of statistical physics of fields for networks [5, 4, 43, 22, 3, 1] permit to quantify the diversity in the traits of the network, with the von Neumann entropy, and the transmission flow of information, with the Helmholtz free energy. As fastness of information flow and sufficiently rich representation of the external world are two competing capabilities, the formation process of a complex networked structure can be interpreted as the macroscopic thermodynamic outcome of the delicate balance between the two, and a variational principle in terms of thermodynamic efficiency can be devised to explain why a certain structure emerge in place of an other [1].
In our work, this variational principle is computationally employed to explain the structural properties of the network subclass composed of sparse heterogeneous hierarchical trees produced by the Portegys protocol that, similar to living systems, need to learn the diversity of the environment they are in contact with and consequently self-organize their structure. The tree-formation undergoes two phase transitions at two different scales: a meso-scale with the emergence of a approximately scale-invariant structural behaviour that progressively involves a larger scales, and a large-scale transition from the irregular scale-invariant behaviour to the regular-scale invariant 1d lattice. Given that the original purpose of the algorithm was an efficient pattern-searching method, we have investigated the pattern-matching efficiency and its positioning on the phase-plane for four different datasets with different level of complexity. When the dataset was more difficult, an optimal efficiency to the pattern-searching task is found to be close to the transition to scale-invariant in the parameter space. Moreover, empirical evidence suggests that the thermodynamic efficiency and the pattern-matching efficiency show some consistency, namely when we look at the global maxima of the two in the $r$-range. Even though the two processes involved possess significant differences, namely the Portegys search phase is a directed process starting from the root node, while the diffusion propagator describes an undirected process propagating from any possible node, we think that is reasonable to expect some similar behavior of the two efficiencies, as both quantities, in the end, estimate a trade-off between the flow of a process and the encoding capabilities of the system in its network structural traits when confronted to the disorder embedded in a dataset of real-world patterns.
In both cases, the "flow" is quantified by a monotonic decreasing quantity as a function of $r$. For the Portegys we have the logarithmic number of computed distances, while for the Laplacian process, we have $\delta F=F+\frac{\log N}{\tau}$, where $F=-\frac{\log Z}{\tau}$. It is easy to convince ourselves that these two quantities describe the flow in the network as the former counts the number of visited nodes across the traversing of the network by the Portegys search process, while the latter is minus the logarithm of $Z$, with $Z$ proportional to the average return probability after time $\tau$ of a continuous time edge-centric random walker [4]. They are not exactly alike as one indeed involves a directed process while the former involves an undirected process; still, they can be used as estimates of the "flow".
We have discussed the denominator of the two efficiencies: $\eta=U/W$ and $\theta=\text{Accuracy}/\log N_{dist}$. We now discuss the numerator. For what concerns the Portegys algorithm, we have observed that there are more descriptive network structures per se as evidenced by the fact that the accuracy is not necessarily a monotonic decreasing function of the control parameter $r$, but it can exhibit higher values at specific $r$-values as illustrated by the cases of the CIFAR10 and the FashionMNIST.
The matter is what makes a certain structure better suited in terms of accuracy to a given task, or equivalently in terms of a complex input-output mapping. Analogously to artificial neural networks, enlarging the number of parameters eventually results in achieving more complicated functions, even though this doesn’t imply a one to one relation between the two. In the case of a network structure the parameters are the network traits themselves, which in the Fourier space become the “network frequencies” $\lambda_{i}$, i.e. the Laplacian eigenvalues. We can measure an effective number of parameters by computing the thermal average of the frequencies contained in the network, precisely the numerator in Equation 5. The internal energy $U$ is the Gibbs average of the Laplacian eigenvalues of the network at the operational resolution scale $\tau$, tantamount to the inverse temperature $\beta$ in statistical physics. Thus, the Gibbs average frequency content is quantified by $U_{\tau}=\operatorname{Tr}[\hat{\rho}\hat{L}]=<\lambda>_{\tau}$ the internal energy, where for scale-invariant networks $<\lambda>_{\tau}=c_{0}/\tau$ [3]. The difference to the case of the artificial neural network is that we can’t directly increase the number of parameters, given a fixed number of samples, but we can let the structure self-organize given an external control parameter $r$.
@@
Notably, we have computed the thermodynamic efficiency at a specific resolution scale $\tau$. We have used the specific heat $C=-dS/d\log\tau$ to identify the $\tau$-resolution which resolves the transition from the trivial structure, that is a single specific heat peak, to the emergence of more complex structures, where the specific heat shows some plateau reflecting the growth of a scale-invariant region where information is integrated at a constant rate. Such $\tau$-resolution scale is akin to an "ultra-violet" cut-off delimiting the relevant frequencies of the network structure across the transition to scale-invariance, i.e. from star graph to amorphous. The operational $\tau$-scale employed roughly identifies modules in the network as at the $r$-bifurcation point we have observed a phase transition when employing the modularity as an order parameter. While the transition point to scale-invariant permits the identification of the network resolution scale where the thermodynamic efficiency needs to be measured, the corresponding transition value of the control parameter $r$ does not coincide with the global maximum in efficiency. We explain this with the statement that a system benefits from the vicinity to a phase transition only when it confronts environments of higher complexity, while it does not benefit from it when the environmental complexity is lower. This is elucidated by the use of four different datasets of increasing complexity and supported by other works in the literature [37, 10, 38].
In terms of practical utility, the methods presented could in principle provide diagnostics to tree-based approaches to classification such as hierarchical clustering, even though a limitation of this study is the investigation of solely the Portegys algorithm. Given a pre-trained tree network structure on a dataset that is unknown, it is possible to qualitatively identify the complexity of the dataset and the ability of the tree structure to cope with that complexity based on the position of the thermodynamic efficiency on the phase portrait without performing any test phase. As discussed in this manuscript, if the trained network is “struggling” in the representation of the dataset, it posits itself close to the phase transition boundary at the cost of a slower pattern search. Interestingly, our analysis on hierarchical tree networks could be also applied, if further developed, to other topics in artificial intelligence. We find it particularly suited to the context of dendritic learning [44], where input units are connected to output units via loop-less paths, to guide the design of the tree-like dendritic structures for biologically inspired computing. In the context of pure neuroscience, it would also be interesting to design experiments where the dendritic structure is related to the capability of the neuron to deal with complicated input-output mapping of signals. Moreover, potential applications include the topic of neuromorphic computing [45, 46, 47] currently looking for a possible solution to the ever-increasing computational power demand of recent years. In the neuromorphic paradigm computing devices are designed such that computation and memory are colocalized on the physical substrate in the attempt to emulate the brain learning capabilities in materio. The full potential of biological neural systems is likely achieved because of the interplay between their complex topological structures and the functional dynamics of a very large number of elements evolving on it [48]. In materio tree-like structures are currently physically implemented. We mention tree structures produced at the critical point of a percolation transition [49, 50, 51] or the electrochemically formed dendrite morphology at the origin of the resistive-switching of some memristive devices [49, 52], thus, their capabilities could be investigated and enhanced based on this work and future developments.